OpenStudy (johnnewman):

The lengths of the legs of a right triangle are 25 cm and 60 cm. What is the length of the hypotenuse? A. 42.5 centimeters B. 35 centimeters C. 65 centimeters D. 9.2 centimeters

pooja195 (pooja195):

Here we would use the Pythagorean theorem. \[\huge~\rm~\bf~a^2+b^2=c^2\] c=hypotenuse the order of a and b don't really matter a=25 b=60 \[\huge~\rm~\bf~25^2+60^2=? \]

OpenStudy (johnnewman):

ok

pooja195 (pooja195):

\[\huge~\rm~25 \times 25=? \]

OpenStudy (johnnewman):

625

OpenStudy (kash_thesmartguy):

Then, what is 60*60?

pooja195 (pooja195):

Good \[\huge~\rm~\bf~60 \times 60=? \]

OpenStudy (johnnewman):

3600

pooja195 (pooja195):

3600+625=?

OpenStudy (johnnewman):

and i add 625 and 3600

pooja195 (pooja195):

Yes

OpenStudy (kash_thesmartguy):

Yep, you got your answer!

pooja195 (pooja195):

That's not the answer

OpenStudy (kash_thesmartguy):

Well, yeah, you then got to square root your answer to get the real answer

pooja195 (pooja195):

@Kash_TheSmartGuy it would really help if you slowed down a bit a gave the user time to answer :)

OpenStudy (johnnewman):

65 is the square root

OpenStudy (kash_thesmartguy):

Ok

pooja195 (pooja195):

c^2=4225 c=65 correct :)

OpenStudy (johnnewman):

sorry i need to slow down a little

pooja195 (pooja195):

Not you i was talking to the other user lol you're fine